OpenStudy (anonymous):

What volume of 1.25M HCl would be required to prepare 180 mL of a 0.500M HCl solution? A. 72 mL B. 0.014 mL C. 2.2 x 103 mL D. 450 mL

OpenStudy (anonymous):

@wingspansmcmemes

OpenStudy (anonymous):

a

OpenStudy (harsha19111999):

M1 * N1 = M2 * N2

OpenStudy (harsha19111999):

x * 1.25 = 180 * 0.5

OpenStudy (harsha19111999):

x = 90/1.25 = 72 ml